OpenStudy (anonymous):

Please help me find this answer!:) Let f(x) = 12/4x+2 Find f(-1). -6 -2 2 6

OpenStudy (anonymous):

@ivettef365

OpenStudy (anonymous):

where you see \(x\) put a \(-1\) \[f(x)=\frac{12}{4x+2}\]\[f(-1)=\frac{12}{4\times (-1)+2}\]